OpenStudy (anonymous):

For how many positive integers n < 100 is (n - 5)(n - 23)(n - 68) positive?

OpenStudy (kinggeorge):

To allow it to be positive, we need one of two cases. Case 1: All three of \((n-5),(n-23),(n-68)\) must be positive. Case 2: Only a single one is positive, and the other two are negative. Let's evaluate these one at a time.

OpenStudy (anonymous):

So It's positive if: \[n > 68 \space or \space 6 < n < 23\]which is 48?

OpenStudy (kinggeorge):

Perfect. Here's the detailed explanation I wrote up. Case 1: \(n\) must be greater than 68, and less than 100. There are exactly \(100-69=31\) possibilities. Case 2: \(5<n<23\) Here, there are exactly \(23-6=17\) possibilities. Add them together, we get\[31+17=48\]choices for \(n\) so that \[(n-5)(n-23)(n-68)>0\]

OpenStudy (anonymous):

Oh ok. So it's very similar to quadratic inequalities right?

OpenStudy (kinggeorge):

Yes, I would call this similar to quadratic inequalities.

OpenStudy (anonymous):

Alright. Thank you!

OpenStudy (kinggeorge):

You're welcome.